a group of ten 6 and 12 volt batteries are wired in a series. The sum of their voltages is 84 volts. HOw many type of each battery is used?
Answer by rfer(10432)

6x+12y=84
x+y=10, x=10-y
6(10-y)+12y=84
60-6y+12y=84
6y=24
y=4
------------------
x+y=10
x+4=10
x=6
--------------
6*6+4*12=84